bob's father is 3 times old as bob. 4 years ago, he was 4 times older. how old is bob?
Found 2 solutions by AAfter Search, ganeshindia:
Answer by AAfter Search(61)

Let the age of Bob be p years.
Then present age of Bob's father is q = 3p
4 years ago, 3p - 4 = 4(p - 4)
=> 3p - 4 = 4p - 16
=> 3p - 4p = 4 - 16
=> -p = -12
p = 12
q = 3 x 12 = 36 years
Hence, present age of Bob is 36 years.

Answer by ganeshindia(11)

Let the present age of bob be x years.
Let the present age of bob's father be y.
Therefore we get, y = 3x.( using the condition "bob's father is 3 times as old as bob).
This gives 3x - y = 0 ..........(1)
Four years ago,
The age of bob = x - 4.
The age of bob's father = y - 4.
Therefore, we get , y - 4 = 4 ( x-4) (using the condition that four years ago he was four times older)
This gives y-4 = 4x - 16
4x - y = 12.........(2)
Solving the linear equations (1) and (2) we get
x = 12 and y=36.

Sol:
Bob's present age x = 12 years
Father's present age y = 36 years